# TANQ: An open domain dataset of table answered questions

### v1

@@ Abstract
Language models, potentially augmented with tool usage such as retrieval are becoming the go-to means of answering questions. Understanding and answering questions in real-world settings often requires retrieving information from different sources, processing and aggregating data to extract insights, and presenting complex findings in form of structured artifacts such as novel tables, charts, or infographics. In this paper, we introduce TANQ, [^1] Datasets at github.com/google-deepmind/tanq the first open domain question answering dataset where the answers require building tables from information across multiple sources. We release the full source attribution for every cell in the resulting table and benchmark state-of-the-art language models in open, oracle, and closed book setups. Our best-performing baseline, GPT4444 reaches an overall F1111 score of $29.1$, lagging behind human performance by $19.7$ points. We analyse baselines’ performance across different dataset attributes such as different skills required for this task, including multi-hop reasoning, math operations, and unit conversions. We further discuss common failures in model-generated answers, suggesting that TANQ is a complex task with many challenges ahead.
@@ 1 Introduction
Understanding and solving problems in real-world scenarios often requires reasoning across multiple documents and data modalities. This includes $(i)$ retrieving information from different sources, $(ii)$ processing and aggregating data to extract insights, and $(iii)$ presenting complex findings in a structured format, for example tables, charts or infographics, to communicate them to readers. Previous studies show that knowledge workers across domains, e.g. finance, science, and economics, spend around $20\%$ of their time searching and gathering information from different files into one document to extract insights and consequently answer information-seeking questions (Chui et al., 2012). As a result, one of most challenging tasks for workers is to aggregate data and turn it into insights. Tables as a structured representation of data are ubiquitous in real-world sources and are commonly used to communicate complex information. Hence, they can be the perfect modality to answer complex questions.
@@
Large language models (LLMs), commonly augmented with external tools such as retrieval, have become the go-to means for answering questions. State-of-the-art question-answering (QA) systems integrate LLMs in various ways, from decomposing complex queries to retrieving documents using external tools or generating context data from knowledge acquired during model training. However, their evaluation is mostly limited to simple datasets, e.g. TabFact (Chen et al., 2020a) or HotpotQA (Yang et al., 2018), whose questions can be answered by reasoning over a single table or text document and generating a short text sequence as answer. This limits the applicability of such systems to perform complex multi-step research explorations. Moreover, it differs from real-world needs where relevant information can be spread across documents and represented in different forms (e.g. text or tables). Moreover, more often than not, generating a short text as answer is not sufficient for more complex information-seeking questions.
@@
We evaluate several state-of-the-art LLMs on TANQ, including close, oracle and open book evaluation setups. Finally, we study model-generated answer tables and discuss common failure cases and challenges related to TANQ. Our evaluation of models across skills can further inform future tools and evaluation setups for LLMs to improve models for complex, information-seeking questions.
To summarise, our contributions are as follows:
We introduce TANQ, the first open-domain question-answering benchmark that requires building answers in form of tables from multiple information sources.
We benchmark state-of-the-art language models in oracle, open and closed book setups, reaching an overall F1111 score of $29.1$ with our best-performing (oracle) baseline.
We evaluate model performance across different dataset characteristics (e.g. reasoning skills and question types) and discuss challenges and common failure types.
Figure 2: TANQ creation pipeline consisting of five steps: 1. Extending QAMPARI questions; 2. Evidence extraction from Wikipedia; 3. Evidence evaluation and answer table extraction; 4. Question rephrasing; 5. Augmentation with additional skills. We include a running example at the bottom with a TANQ example at the end.
@@ 2 Related Work
The work closest to ours with respect to input and output modalities, is MultiTabQA Pal et al. (2023). MultiTabQA seeds on the Spider dataset, a text-to-SQL dataset containing SQL queries, database tables, and natural language translations of the queries. Pal et al. (2023) use table names occurring in SQL queries to extract the input tables and query Spider databases for answer table generation. While the dataset also generates answers as tables, it has certain limitations: $(i)$ the benchmark input is limited to tables; $(ii)$ the input and output tables are highly structured database tables which differ from real-world scenarios where tables occur in documents and websites in various formats; $(iii)$ the questions are limited to SQL queries formulated in natural language.
@@ 1. Simple Which Belgian Grand Prix did Michael Schumacher win, and when did that happen?
Table 2: An overview of question types in TANQ and additional skills we use for augmenting the TANQ questions for more complex reasoning beyond retrieval.
@@ 3.2 Preliminaries
QAMPARI. We use QAMPARI as seed dataset which is an open-domain QA dataset with lists of entities as answers. QAMPARI further includes Wikipedia text as supporting evidence for each entry of the answer list. Different to prior QA datasets with short textual answers, they align to natural questions which require a list of answers extracted from multiple sources. The dataset is semi-automatically created with Wikidata and Wikipedia as data sources. Moreover, Amouyal et al. (2022) ask human annotators to validate each QAMPARI question.
WikiData. Following Amouyal et al. (2022), we use Wikidata (WD) as a source for question generation. Wikidata (Erxleben et al., 2014) is a collaborative knowledge graph consisting of triples of entities and relations, i.e. $(e_{1},r,e_{2})$. Entities $e_{i}$ are values (e.g. $1990$) or items that represent a real-world concept, object, or a topic. [^2] https://en.wikipedia.org/wiki/Wikidata Relations are edges connecting entities, e.g. $(DonaldTrump,instanceOf,human)$. Whereas “Donald Trump” is the head entity and “human” the tail entity. We follow the notion for formal queries over Wikidata introduced by Amouyal et al. (2022) for QAMPARI. Applying a relation $r$ as query over WD items $e_{i}$ results in a set of (tail) entities: $[[r(e)]]=\{e_{i}|(e_{i},r,e)\in\text{WD}\}$.
@@
Type Count Freq (%)
WD Item 57.6k 79.1
Time 11.9k 16.4
Numeric 3.2k 4.4
Text 46 0.1
Table 3: Extension entities ($1$st pipeline step).
@@
1. Simple 504 36.1
2. Composition 320 22.9
3. Intersection 571 40.9
@@
1. No skill 385 27.6
2. Filtering numeric 85 6.1
3. Filtering time 236 16.9
4. Filtering entity 342 24.5
5. Date-to-year conversion 310 22.2
6. Quantity conversion 95 6.8
7. Time calculation 45 3.2
8. Approximation 94 6.7
@@
1. No skill 385 27.6
2. One skill 843 60.4
3. Two skills 137 9.8
4. Three skills 30 2.2
@@ 3.3 TANQ benchmark pipeline
Figure 2 provides an overview of the TANQ pipeline outlining all steps for data collection, processing and evaluation. We use QAMPARI, Wikidata and Wikipedia as data sources, as well as PaLM-2 Anil et al. (2023b) for paraphrasing and validation. Applying a five-step process, we create various dataset entries with different question types and skills.
@@
Next, we collect for each extension triple ($e_{a}$, $r_{ext}$, $e_{ext}$) supporting evidence using Wikipedia as an evidence source. Hereby, we search for supporting text, tables and infoboxes in the WP articles of $e_{a}$ and $e_{ext}$. We apply simple heuristics and search for mentions of $e_{a}$ in the $e_{ext}$ article and vice versa. Moreover, we extend our queries with additional (heuristic-based) query words, for example considering different forms of representing numeric or time-related information occurring in queries.
@@
To evaluate the correctness of the previously collected evidence texts, tables and infoboxes, we employ PaLM-2 as an evidence evaluator. We prompt the LLM to evaluate the extracted evidence in a natural language inference setting. For each extension triple $(e_{a}$, $r_{ext}$, $e_{ext})$, we construct template-based sentences $s$: “$\langle e_{a}\rangle\langle r_{ext}\rangle\langle e_{ext}\rangle$”, e.g. “Hey Ram composer Ilaiyaraaja” for row two in Figure 1. We query the LLM to label the sentence s as “supported”, “refuted”, “not enough information” based on the provided evidence in form of a sentence, table or infobox entry. We then only consider the triples supported by at least one piece of evidence. [^3] See Appendix A for the prompt template.
Figure 3: Baseline models evaluated on TANQ: closed book, oracle, and open book with augmented tools.
Answer table. Each extended relation corresponds to a column in the answer table (e.g. column “Composer” in Figure 1). Since the question in Figure 1 is extended with two additional relations (i.e. composer and release year), the resulting answer table has three columns. Hence, each cell in the answer table corresponds to a WD triple. For example, cell “Ilaiyaraaja” corresponds to triple ((((Hey Ram, composer, Ilaiyaraaja)))).
@@
This step increases the naturalness of template-based extension questions generated in Step 1. Similarly to the earlier step, we prompt a PaLM-2 model in a few-shot setting. To ensure the question meaning is preserved during rephrasing, we add structured annotations in parenthesis to questions with the name of each relation (e.g “Which Indian movies were both directed (directed_by) and written by (written_by) Kamal Haasan?”). We run up to 5 iterations of rephrasing and stop if all relations are present, or discard the question otherwise. [^4] See Appendix A for the prompt template.
@@ 3.4 Dataset statistics and analysis
Overall, the generated TANQ dataset has $1395$ entries, $36.1\%$ of question type simple, $40.9\%$ intersection, and $22.9\%$ composition questions. Approximately $72.4\%$ dataset instances (i.e. $,01$ 010) require at least one additional skill to answer the question. See Table 4 for a breakdown of skills and question types in TANQ. TANQ questions have an average a length of $21$ tokens and require extracting information about three relations on average which are related to the answer entities (similarly to composer and release year in Figure 1). On average, TANQ answer tables have $6.7$ rows and four columns. Table 3 given an overview of extension entities used. We further plan to release a TANQ training set with approximately $42k$ samples.
@@
PaLM-$2$ 16.8 17.7 15.4 13.3 14.1 11.6 20.1 23.8 19.5 17.3 19.4 16.2
Gemini 56.1 15.4 14.3 53.8 10.1 9.7 48.9 16.2 13.9 52.6 14.5 13.1
GPT4 44.3 29.7 29.0 34.8 20.1 22.7 35.7 23.8 22.0 45.7 29.4 29.1
Human 54.6 48.6 48.8
@@
PaLM-$2$ 17.2 17.1 13.6 17.0 10.3 10.4 26.4 22.0 21.1 20.9 17.5 15.9
Gemini 32.8 11.0 8.4 37.1 6.4 5.9 31.8 17.9 16.0 33.4 12.7 10.9
GPT4 32.5 20.7 17.4 27.3 12.5 12.6 45.1 26.0 24.3 37.3 21.5 19.6
@@
Tool LM 13.4 12.5 9.7 21.5 14.1 15.1 19.4 19.2 15.7 17.3 15.8 13.2
Table 5: Baseline performance by question type. For all question types, we observe GPT4444 to outperform other baselines in oracle, closed, and open book settings, resulting in an overall F1111 score of $29.1$ for all question types.
@@
PaLM-$2$ 16.7 18.0 16.0 16.2 16.0 14.1 23.0 16.7 15.2 30.2 25.2 23.6 19.9 17.3 16.4 12.9 6.4 7.3 15.4 12.0 12.0
Gemini 28.2 14.5 11.1 23.7 13.9 11.3 37.3 10.2 8.2 36.0 18.0 16.0 19.0 15.0 13.6 51.5 1.6 1.8 33.0 9.8 9.3
GPT4 36.7 16.6 15.1 34.7 25.8 22.1 31.5 23.0 22.4 49.7 22.0 20.2 71.6 7.3 5.7 44.6 1.4 2.0 36.5 19.1 19.3
@@
Tool LM 16.7 21.1 17.1 15.7 11.1 10.9 21.0 20.6 16.4 20.3 17.5 16.0 9.7 10.8 9.7 28.9 13.0 10.7 16.0 19.9 15.2
Table 6: Baseline performance by skills: Filtering with numerical/datetime/entity conditions, date-to-year (Date2Year), quantity conversion, time calculation, and approximation. In the oracle setting GPT4444 outperforms other baslines for most skills, while PaLM-$2$ performs better on some skills in closed book evaluation.

### v3

@@ Abstract
Language models, potentially augmented with tool usage such as retrieval are becoming the go-to means of answering questions. Understanding and answering questions in real-world settings often requires retrieving information from different sources, processing and aggregating data to extract insights, and presenting complex findings in form of structured artifacts such as novel tables, charts, or infographics. In this paper, we introduce TANQ, [^1] Dataset’s at github.com/google-deepmind/tanq the first open domain question answering dataset where the answers require building tables from information across multiple sources. We release the full source attribution for every cell in the resulting table and benchmark state-of-the-art language models in open, oracle, and closed book setups. Our best-performing baseline, Gemini Flash reaches an overall F1111 score of $60.7$, lagging behind human performance by $12.3$ points. We analyse baselines’ performance across different dataset attributes such as different skills required for this task, including multi-hop reasoning, math operations, and unit conversions. We further discuss common failures in model-generated answers, suggesting that TANQ is a complex task with many challenges ahead.
@@ 1 Introduction
Understanding and solving problems in real-world scenarios often requires reasoning across multiple documents and data modalities. This includes $(i)$ retrieving information from different sources, $(ii)$ processing and aggregating data to extract insights, and $(iii)$ presenting complex findings in a structured format, for example a table or infographics, to communicate them to readers.
@@
Previous studies show that knowledge workers across domains, e.g. finance, science, and economics, spend around $20\%$ of their time searching and gathering information from different files into one document to extract insights and consequently answer information-seeking questions (Chui et al., 2012). As a result, one of most challenging tasks for workers is to aggregate data and turn it into insights. Tables as a structured representation of data are ubiquitous in real-world sources and are commonly used to communicate complex information. Hence, they can be the perfect modality to answer complex questions.
@@
Large language models (LLMs), often enhanced with external tools, have become a primary method for various application such as answering questions. State-of-the-art question-answering (QA) systems integrate LLMs in various ways, from decomposing complex queries to retrieving documents using external tools or generating context data from knowledge acquired during model training. However, their evaluation is mostly limited to simple datasets, e.g. TabFact (Chen et al., 2020a) or HotpotQA (Yang et al., 2018), whose questions can be answered by reasoning over a single table or text document and generating a short text sequence as answer. This limits the applicability of such systems to perform complex multi-step research explorations. Moreover, it differs from real-world needs where relevant information can be spread across documents and represented in different forms (e.g. text or tables). More often than not, generating a short text as answer is not sufficient for complex information-seeking questions.
@@
We evaluate several state-of-the-art LLMs on TANQ, including close, oracle and open book evaluation settings. Finally, we study model-generated answer tables and discuss common failure cases and challenges related to TANQ. Our evaluation of models across skills can further inform future tools and evaluation setups for LLMs to improve models for complex, information-seeking questions.
Our contributions are as follows:
(a) We introduce TANQ, the first open-domain question-answering benchmark that requires building answers in form of tables from multiple information sources.
(b) We benchmark state-of-the-art language models in oracle, open and closed book setups, reaching an overall F1111 score of $60.7$ with our best-performing (oracle) baseline.
(c) We evaluate model performance across different dataset characteristics, discuss challenges and common failure types.
Figure 2: TANQ creation pipeline consisting of five steps: 1. Extending QAMPARI questions with additional relations based on Wikidata; 2. Evidence extraction (text, tables, infoboxes) from Wikipedia articles; 3. Evidence evaluation and (gold) answer table generation; 4. Rephrasing of question from first step; 5. Augmentation with additional skills to generate complex. We include a running example at the bottom.
@@ 2 Related Work
The work closest to ours with respect to input and output modalities, is MultiTabQA Pal et al. (2023). MultiTabQA seeds on the Spider dataset, a text-to-SQL dataset containing SQL queries, database tables, and natural language translations of the queries. Pal et al. (2023) use table names occurring in SQL queries to extract the input tables and query Spider databases for answer table generation. While the dataset also generates answers as tables, it has certain limitations: $(i)$ the benchmark input is limited to tables; $(ii)$ the input and output tables are highly structured database tables which differ from real-world scenarios where tables occur in documents and websites in various formats; $(iii)$ the questions are limited to SQL-based queries.
@@ 1. Simple Which Belgian Grand Prix did Michael Schumacher win, and when did that happen?
Table 2: An overview of TANQ question types and skills we use for augmenting the questions for more complex reasoning in the final step of the TANQ pipeline in Figure 2.
@@ 3.2 Preliminaries
QAMPARI. We use QAMPARI as seed dataset which is an open-domain QA dataset with lists of entities as answers (Amouyal et al., 2022). QAMPARI further includes Wikipedia text as supporting evidence for each entry of the answer list. Different to prior QA datasets with short textual answers, they align to natural questions which require a list of answers extracted from multiple sources. The dataset is semi-automatically created with Wikidata and Wikipedia as data sources and evaluated by human annotators.
WikiData. Following Amouyal et al. (2022), we use Wikidata (WD) as a source for question generation. Wikidata (Erxleben et al., 2014) is a collaborative knowledge graph consisting of triples of entities and relations, i.e. $(e_{1},r,e_{2})$. Entities $e_{i}$ are values (e.g. $1990$) or items that represent a real-world concept, object, or a topic. [^2] https://en.wikipedia.org/wiki/Wikidata Relations are edges connecting entities, e.g. (DonaldTrump, instanceOf, human). Whereas Donald Trump is the head entity and human the tail entity (Krishna et al., 2022). We follow the notion for formal queries over Wikidata introduced by Amouyal et al. (2022) for QAMPARI. Applying a relation $r$ as query over WD items $e_{i}$ results in a set of (tail) entities: $[[r(e)]]=\{e_{i}|(e_{i},r,e)\in\text{WD}\}$.
@@
Type # % Type # %
WD Item 57.6k 79.1 Numeric 3.2k 4.4
Time 11.9k 16.4 Text 46 0.1
Table 3: Types of Wikidata entities we use in the $1$st pipeline step (see Figure 2) to extend QAMPARI and generate TANQ questions. We mostly use Wikidata (WD) items.
@@
1. Simple 428 39.9
2. Composition 164 15.3
3. Intersection 482 44.9
@@
1. No skill 160 14.9
2. Filtering numeric 66 6.1
3. Filtering time 190 17.7
4. Filtering entity 268 25.0
5. Date-to-year conversion 230 21.4
6. Quantity conversion 64 6.0
7. Time calculation 23 2.1
8. Approximation 73 6.8
@@
1. No skill 300 27.9
2. One skill 657 61.2
3. Two skills 94 8.8
4. Three skills 23 2.1
@@ 3.3 TANQ benchmark pipeline
Figure 2 provides an overview of the TANQ pipeline outlining all steps for data collection, processing and evaluation. We use QAMPARI, Wikidata and Wikipedia, as well as PaLM-2 Anil et al. (2023b) for paraphrasing and validation. [^3] Prompts provided in the appendix (Figure 4 and 6).
@@
Next, we collect for each extension triple ($e_{a}$, $r_{ext}$, $e_{ext}$) supporting evidence using Wikipedia as an evidence source. Hereby, we search for supporting text, tables and infoboxes in the WP articles of $e_{a}$ and $e_{ext}$. We apply simple heuristics and search for mentions of $e_{a}$ in the $e_{ext}$ article and vice versa. Moreover, we extend our queries with additional (heuristic-based) query words, for example considering different formats of how numbers and dates are represented in queries.
@@
To evaluate the correctness of the previously collected evidence texts, tables and infoboxes, we employ PaLM-2 as an evidence evaluator. We prompt the LLM to evaluate the extracted evidence in a natural language inference setting. For each extension triple $(e_{a}$, $r_{ext}$, $e_{ext})$, we construct template-based sentences $s$: “$\langle e_{a}\rangle\langle r_{ext}\rangle\langle e_{ext}\rangle$”, e.g. “Hey Ram composer Ilaiyaraaja” for row two in Figure 1. We query the LLM to label the sentence s as “supported”, “refuted”, “not enough information” based on the provided evidence in form of a sentence, table or infobox entry. We then only consider the triples supported by at least one piece of evidence. [^4] See Figure 4 for exact prompt we used.
Figure 3: Prompts for baseline evaluation on TANQ: closed book, oracle, and open book with augmented tools.
Answer table. Each extended relation corresponds to a column in the answer table (e.g. column “Composer” in Figure 1). Since the question in Figure 1 is extended with two additional relations (i.e. composer and release year), the resulting answer table has three columns. Hence, each cell in the answer table corresponds to a WD triple. For example, (Hey Ram, composer, Ilaiyaraaja) for cell “Ilaiyaraaja”. Some generated answer tables contained multiple entries in a single cell, as seen in the genre column of Table 7. To generate realistic tables, we filtered out samples with more than five entries in any single cell, resulting in a test set of 1,074 TANQ samples for evaluation.
@@
This step increases the naturalness of template-based extension questions generated in Step 1. Similarly to the earlier step, we prompt a PaLM-2 model in a few-shot setting. To ensure the question meaning is preserved during rephrasing, we add structured annotations in parenthesis to questions with the name of each relation (e.g “Which Indian movies were both directed (directed_by) and written by (written_by) Kamal Haasan?”). We run up to 5 iterations of rephrasing and stop if all relations are present, or discard the question otherwise. [^5] See Figure 6 for exact prompt we used.
@@ 3.4 Dataset statistics and analysis
Overall, the generated TANQ dataset has $1395$ entries, $36.1\%$ of question type simple, $40.9\%$ intersection, and $22.9\%$ composition questions. We further plan to release a TANQ training set with approximately $42k$ samples. Approximately $72.4\%$ dataset instances (i.e. $,01$ 010) require at least one additional skill to answer the question. See Table 4 for a breakdown of skills and question types in TANQ. TANQ questions have an average a length of $21$ tokens and require extracting information about three relations on average. On average, TANQ answer tables have $6.7$ rows and four columns. See Table 3 for WD entity types used for extending QAMPARI questions.
3.5 Manual evaluation of the pipeline
We manually evaluated $100$ TANQ samples to assess if noise introduced through the dataset’s automated generation pipeline. The focus of the analysis was twofold. First, examining the initial questions sourced from the QAMPARI dataset and, second, evaluating the generation of TANQ questions over the different pipeline steps. We identified five issues which are discussed below.
Propagation of QAMPARI issues.
While occurring only for a small subset of questions, one observation was that some errors present in the initial QAMPARI questions were propagated through the pipeline without being corrected. In $2$ out of the $100$ samples, the original QAMPARI questions contained issues, one was carried into TANQ while the other was fixed. Examples include: “Robert Benton is the screenwriter and director of which software?” This question confuses a film with software as we see by considering the requested attributes, i.e., screenwriter and director.
While these logical errors were carried forward into TANQ grammatical errors, for example, were fixed during rephrasing with LLMs: “Black Sabbath had who as a member?” — issues phrasing that was later improved to “Who were the members of Black Sabbath […]?”.
Partially revealed answers.
In three out of hundred evaluted questions, the question itself partially contained the answer, rendering the query less meaningful: “Eon Productions and Harry Saltzman produced what series and what was their genre, director of photography, duration in second, producer, country of origin, composer, publication date, narrative location, screenwriter, and director?” — in this case, the “producer” is already mentioned in the question.
Grammar and expression errors.
Minor grammatical issues introduced through the TANQ pipeline, were found in $2$ out of the $100$ samples. In one cases, this was fixed during the rephrasing process, while other remained unresolved. For example, “+89 minute” was rephrased to “89 minutes” in the sentence, “Which film was directed and produced by Mel Brooks and what was their composer, genre, duration? Filter the answer table for duration equal to or larger than +89 minute.” However, “Larger than 89 minutes” remained, but “longer than 89 minutes” would be more appropriate for this sentence.
Ordering of relations.
While our evaluation metrics is not sensitive to the ordering of columns within the answer table, the order of relation appeared confusing for some questions. This occurred for seven out of the evaluated $100$ examples as shown below: “What filmmaker directed a movie written by Val Guest and what was their date of birth, place of death, date of death, and place of birth?” — the order of personal attributes (i.e., date and place of birth, date and place of death) seems unnatural as date of birth would naturally be followed by place of birth.
Ambiguous or unclear relations.
In four out of the $100$ samples, certain relations were ambiguous or unclear, causing confusion about what exactly the question was asking for. Such as, “Who directed the movie which had the screenwriting done by Marc Norman and what was their date of birth, language spoken, written or signed, place of birth, sex or gender, country of citizenship, occupation?” — here, the attribute “language spoken, written or signed” is not clearly related to the director in question.
While we observed five different issues during the manual evaluation of the TANQ pipeline, the occurred only in a limited number of samples out of $100$ evaluated one. Thus, while these errors are present, they are not widespread enough to strongly impact the overall quality and utility of the TANQ dataset.
@@ 3.5 Manual evaluation of the pipeline
Gemini Pro 41.5 38.1 37.6 39.8 36.7 36.6 42.4 40.6 40.0 41.5 38.9 38.4
Gemini Flash 66.4 59.0 60.4 67.4 58.4 60.4 66.8 59.3 61.1 66.8 59.0 60.7
[…]
Human 81.2 74.0 73.0
@@
PaLM-$2$ 55.5 48.1 49.6 50.6 45.7 46.6 52.1 45.5 46.6 52.9 46.4 47.6
Gemma 38.9 26.7 27.9 37.4 25.8 26.6 39.3 26.2 27.7 38.8 26.3 27.5
Gemini Pro 46.9 30.3 31.1 43.8 29.5 30.4 47.8 32.0 33.5 46.6 30.9 32.0
Gemini Flash 37.4 29.0 31.3 36.2 27.2 29.6 36.6 28.8 30.7 36.8 28.5 30.7
[…]
@@
Tool LM 24.2 20.3 18.4 43.6 33.0 36.8 50.8 44.2 46.6 39.7 33.4 34.5
Table 5: Baseline performance by question type. For all question types, we observe Gemini Flash ($60.7$ F1111) and PaLM-$2$ ($47.6$ F1111) to outperform other baselines in oracle and closed book setting respectively, lagging $12.3$ and $25.4$ points behind the human baseline of $73.0$.
@@
PaLM-$2$ 50.5 46.3 46.9 53.9 47.1 48.0 54.9 48.5 49.7 51.4 44.7 46.1 51.5 45.2 46.8 53.1 48.3 50.1 52.3 43.4 45.7
Gemma 37.4 26.5 27.8 37.1 26.4 27.3 40.3 26.2 27.6 40.4 27.4 28.7 37.4 23.8 25.2 39.2 27.4 29.3 40.8 27.5 29.6
Gemini Pro 50.2 28.3 29.3 48.5 31.4 32.4 46.3 31.2 32.4 46.1 30.6 31.9 48.0 31.4 33.5 51.4 33.9 35.3 42.8 30.2 31.4
Gemini Flash 38.7 29.5 32.1 35.8 26.7 29.1 36.2 27.7 29.9 35.9 27.8 29.7 36.6 28.0 29.1 34.8 27.5 30.1 37.1 27.1 30.0
[…]
@@
Tool LM 48.7 39.1 43.1 36.1 29.7 30.3 37.6 33.0 34.2 53.1 37.8 41.9 33.8 28.4 30.2 25.8 18.2 21.1 48.3 44.0 45.9
Table 6: Baseline performance by skills required to answer the question: Filtering with numerical/datetime/entity conditions, date-to-year (Date2Year), quantity conversion, time calculation, and approximation. In the oracle setting Gemini Flash performs best across all skills, while PaLM-$2$ performs better in the closed book evaluation.